how do i solve this as a single logarithm
log7 - 1og9

Answer by Earlsdon(6294)

Remember the quotient rule for condensing logarithms:
Log(A) - Log(B) = Log(A/B)
For your problem, this would translate into:
Log(7) - Log(9) = Log(7/9) = -0.1091444...